OpenStudy (anonymous):

solve: lnx + ln(x-1) = 1

zepdrix (zepdrix):

\[\Large \ln x+\ln(x-1)=1\] First we want to apply this rule of logarithms:\[\large \color{royalblue}{\log(a)+\log(b)=\log\left(a\cdot b\right)}\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so it would be ln(x^2+x)

zepdrix (zepdrix):

\[\Large \ln(x^2+x)=1\] Mmm ok looks good. So for the next step, we can either exponentiate each side, which will deal with the logarithm. Or, if you remember how to rewrite a log in exponent form we can skip that step.

OpenStudy (anonymous):

so it would be \[e ^1 = x^2 +x\]

zepdrix (zepdrix):

Yes good! :)

OpenStudy (anonymous):

but then I have to solve for x which confuses me

zepdrix (zepdrix):

Hmm I guess we could start by subtracting e from each side,\[\Large 0=x^2+x-e\]And next let's throw this into the Quadratic Formula to find solutions for x,\[\Large x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

wouldnt it be like a weird answer, like (x-ex)(x+ex) or something like that

zepdrix (zepdrix):

Ya it's definitely going to be a weird answer :) It won't factor nicely so we have to use the quadratic formula.

zepdrix (zepdrix):

\[\Large x=\frac{-b\pm\sqrt{b^2-4ac}}{2a} \quad\to\quad x=\frac{-1\pm\sqrt{1^2-4\cdot1\cdot(-e)}}{2\cdot1}\] Understand how I plugged those in? :o

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so it would be (-1 +- root3e)/2 ?

zepdrix (zepdrix):

I think it's going to be, \[\Large x=\frac{-1\pm\sqrt{1+4e}}{2}\]

zepdrix (zepdrix):

But we'll need to do a little bit of extra work. We have two possible solutions for x. We need to see if both of them will work for our `original function`. Let's punch these into the calculator, it will be a lot easier to work with that way.\[\large x=\frac{-1-\sqrt{1+4e}}{2}\quad=\quad?\]\[\large x=\frac{-1+\sqrt{1+4e}}{2}\quad=\quad?\]

OpenStudy (anonymous):

approx -2.22 and 1.22

zepdrix (zepdrix):

Mmmm ok good good, that looks right. Will either of those x values give us a problem if we try to plug it into our function?\[\Large \ln x+\ln(x-1)=1\]If you're not sure, try putting them in a see what happens.\[\Large \ln(-2.22) \quad=\quad ?\]

OpenStudy (anonymous):

it says error on my calculator :p

zepdrix (zepdrix):

Interesting! :O

zepdrix (zepdrix):

Ya we can't plug negative numbers into a log :) Try to remember that!! The domain of ln(x) is \(\large (0,\infty)\). So we can't use the x=-2.22 solution.

zepdrix (zepdrix):

So our final answer would be,\[\Large x=1.22 \qquad\text{or}\qquad x=\frac{1+\sqrt{1+4e}}{2}\] Depending on whether you prefer to leave it exact or decimal form :)

zepdrix (zepdrix):

Woops I typoed,\[\Large x=\frac{-1+\sqrt{1+4e}}{2}\]

OpenStudy (anonymous):

so there is one x answer and that will be my domain? so it's (-infinity,1)u(1.22,infinity) ?

zepdrix (zepdrix):

There is one x value that solves the problem. That's the value we found. Do you also need to specific the domain of the function?

OpenStudy (anonymous):

no, i didn't need a domain. I thought I did :p thanks so much! you're a life saver!

zepdrix (zepdrix):

yay team \c:/